# Spatially-guided Temporal Aggregation for Robust Event-RGB Optical Flow Estimation

### v1

@@ VI Conclusion and limitation
This paper explores the fusion of event and frame data for optical flow estimation, leveraging the complementary strengths of both modalities. Our approach integrates the spatial stability of frames with the high temporal resolution of events, introducing a novel paradigm for robust optical flow estimation. Specifically, we reorganize the two modalities into a spatially robust guiding modality, extracting stable features to guide the aggregation of temporally-dense event features. The aggregation module leverages transformer architectures, where guided features are used to enrich event features and guide the temporal aggregation. The proposed guidance strategy enhances cross-modal interactions and fully exploits the complementary strengths between events and frames, enabling more effective fusion. Additionally, we analyze the impact of different contexts in optical flow estimation and introduce a compact spatiotemporal context to replace single-modality contexts. Experimental results demonstrate that our fusion strategy achieves rapid convergence and state-of-the-art accuracy with minimal iterations. Comprehensive ablation studies further validate the robustness and effectiveness of our method.

### v2

@@ VI Conclusion
This paper explores the fusion of event and frame data for optical flow estimation, leveraging the complementary strengths of both modalities. Our approach integrates the spatial stability of frames with the high temporal resolution of events, introducing a novel paradigm for robust optical flow estimation. Specifically, we reorganize the two modalities into a spatially robust guiding modality. The aggregation module leverages transformer architectures, where guided features are used to enrich event features and guide the temporal aggregation. The proposed strategy enhances cross-modal interactions and fully exploits the complementary strengths between events and frames. Additionally, we introduce a compact spatiotemporal context to replace single-modality contexts. Experimental results demonstrate that our fusion strategy achieves rapid convergence and state-of-the-art accuracy with minimal iterations. Comprehensive ablation studies further validate the robustness and effectiveness of our method.